OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x-4y=6 3x+y=-8 A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3)

OpenStudy (anonymous):

The answer should be C. (-2,-2) plug in the x value for all x's in the equation and y value for all y's in the equation (-2)-4(-2)=6 -2+8=6 3(-2)+(-2)=-8 -6-2=-8